# Face-Human-Bench: A Comprehensive Benchmark of Face and Human Understanding for Multi-modal Assistants

### v2

@@ Abstract
Faces and humans are crucial elements in social interaction and are widely included in everyday photos and videos. Therefore, a deep understanding of faces and humans will enable multi-modal assistants to achieve improved response quality and broadened application scope. Currently, the multi-modal assistant community lacks a comprehensive and scientific evaluation of face and human understanding abilities. In this paper, we first propose a hierarchical ability taxonomy that includes three levels of abilities. Then, based on this taxonomy, we collect images and annotations from publicly available datasets in the face and human community and build a semi-automatic data pipeline to produce problems for the new benchmark. Finally, the obtained Face-Human-Bench comprises a development set with 900 problems and a test set with 1800 problems, supporting both English and Chinese. We conduct evaluations over 25 mainstream multi-modal large language models (MLLMs) with our Face-Human-Bench, focusing on the correlation between abilities, the impact of the relative position of targets on performance, and the impact of Chain of Thought (CoT) prompting on performance. Moreover, inspired by multi-modal agents, we also explore which abilities of MLLMs need to be supplemented by specialist models.
@@ 1 Introduction
Faces and humans are always the most crucial elements of photos and videos in our everyday lives. Consequently, they are also critical focuses in multi-modal AI applications. In the past two years, ChatGPT (OpenAI, 2023a) and GPT-4 (OpenAI, 2023b) have achieved great success with impressive instruction-following and multi-modal understanding capabilities respectively. Numerous excellent works (Liu et al., 2023b; Zhu et al., 2024; Dai et al., 2023; Bai et al., 2023) from the open-source community have followed, collectively presenting the immense potential of multi-modal assistants. Since faces and humans are central to social interaction, a deep understanding of this information can make multi-modal assistants achieve improved response quality and broadened application scope. For instance, in movie understanding (Yue et al., 2023; Han et al., 2023; Wang et al., 2024), identifying characters is a prerequisite for multi-modal assistants to describe the plot accurately. In multi-modal human-computer interaction (Fu et al., 2024), perceiving expressions and body language can help multi-modal assistants accurately understand the context, generating more personalized and humanized responses. In media forensics (Liu et al., 2024b; c; Jia et al., 2024), determining whether deepfake artifacts exist on a face is crucial for multi-modal assistants to detect misinformation.
Comprehensive and scientific evaluation is the foundation for researching applications of multi-modal assistants related to “faces and humans.” Existing benchmarks Fu et al. (2023); Li et al. (2023a); Liu et al. (2023c) for large multi-modal models typically involve limited abilities of face and human understanding, such as celebrity recognition, action recognition, identity reasoning, and social relation, leaving many important abilities unexplored. On the other hand, since face and human understanding is one of the earliest research topics in artificial intelligence, there are numerous datasets available for evaluating the performance of specialist models. The images and annotations from these datasets can serve as original material to evaluate multi-modal assistants.
As the starting point of our evaluation, we propose a hierarchical ability taxonomy, as shown in Figure 1. This taxonomy consists of three levels. Level-1 (L1) has two perspectives to study: from the target perspective, L1 includes face understanding and human understanding; from the cognitive process perspective, L1 includes perception and reasoning. Subsequently, we incorporate finer-grained abilities into the taxonomy and categorize them into 10 Level-2 (L2) and 18 Level-3 (L3) ability dimensions. Then, based on this taxonomy, we collect datasets from the face and human community and use a semi-automatic data pipeline to transform original images and annotations into multi-modal QAs. The final obtained benchmark called Face-Human-Bench, including a development set with 900 problems and a test set with 1800 problems, supporting evaluations in both English and Chinese. For ease of evaluation, we adopt multiple-choice as the problem format following MMBench (Liu et al., 2023c) and SEED-Bench (Li et al., 2023a).
@@
Figure 1: The hierarchical ability taxonomy for evaluating face and human understanding abilities. We construct the Face-Human-Bench based on this taxonomy. The proportion of the sectors represents the weight of the corresponding abilities in the overall score on the Face-Human-Bench.
In the literature, multi-modal assistants can be broadly categorized into two types: (1) Multi-modal large language models (MLLMs), which achieve end-to-end output by aligning visual information to the language domain with visual instruction-tuning (Liu et al., 2023b). (2) Multi-modal agents (Wu et al., 2023; Yang et al., 2023), where LLMs decide when to call specialist models to solve particular problems and then integrate the outputs of these specialist models. Compared to multi-modal agents, MLLMs generally have better multi-modal perception and reasoning abilities with more effective relationship modeling across modalities. In this study, the first research question (RQ1) is: “How do existing MLLMs perform in face and human understanding?” In this question, we focus on (a) the performance of 25 mainstream MLLMs, (b) the correlation between abilities at different levels, (c) the impact of the relative position of targets on performance, and (d) the impact of Chain of Thought (CoT) prompting on performance. Meanwhile, for face and human understanding tasks in which specialist models significantly outperform MLLMs, we can draw inspiration from multi-modal agents by utilizing the output of these specialist models to enhance the responses of multi-modal assistants. Thus, the second research question emerges (RQ2): In the field of face and human understanding, which tasks’ specialist models can achieve significantly better performance than current MLLMs?
@@
In response to RQ1, our main findings are as follows: (a) The Face-Human-Bench effectively distinguishes the abilities of MLLMs in faces and human understanding. Under the zero-shot setting, the best-performing closed-source model, GPT-4o (OpenAI, 2024), does not perform as well as the best open-source model, InternVL-Chat-v1.2-Plus (Chen et al., 2023). (b) The correlation coefficients can reveal correlations between abilities at different levels. At L2 and L3, there are some ability groups in which the ability dimensions exhibit significant positive correlations between each pair. (c) Many models show substantial performance differences on the same task with different relative positions of targets. We design a new metric called the relative position sensitivity score (RPSS) to measure this phenomenon. On this metric, InternLM-XComposer2-VL-7B (Dong et al., 2024) performs the best, indicating that its performance is almost unaffected by the relative position of targets. (d) Introducing hints and CoT instructions into the prompts significantly improves the performance of the closed-source model GPT-4o, but has no effect on the open-source model, InternVL-Chat-v1.2-Plus. In response to RQ2, we find that in tasks of deepfake detection, crowd counting, and face recognition (under challenging scenarios), the performance of MLLMs is significantly inferior to that of corresponding specialist models. Therefore, we recommend incorporating specialist models in applications requiring these abilities to help improve the response quality of multi-modal assistants.
@@
  • We propose the Face-Human-Bench, the first benchmark dedicated to evaluating multimodal assistants’ face and human understanding abilities. The Face-Human-Bench is based on a three-level ability taxonomy and supports both English and Chinese.
@@
  • We explore which specialist models significantly outperform MLLMs in certain face and human understanding tasks. Based on this, we provide suggestions for enhancing the response quality of multi-modal assistants.
@@ 2.1 Hierarchical Ability Taxonomy
As shown in Figure 1, the proposed ability taxonomy includes three levels. Level-1 (L1) has two research perspectives. From the target perspective, L1 includes face understanding and human understanding. From the cognitive process perspective, L1 includes perception and reasoning. In our evaluation, perception involves direct comprehension of only one target, while reasoning requires synthesizing information from multiple targets and environments to conclude. There are ten abilities in total at Level-2 (L2). Five are focused on faces: facial attribute recognition, age estimation, facial expression recognition, face attack detection, and face recognition, and five are focused on humans: human attribute recognition, action recognition, spatial relation understanding, social relation understanding, and person re-identification. It should be noted that at L2, there are 6 abilities under perception and 4 abilities under reasoning. Level-3 (L3) further refines the ability dimensions at L2. Facial expression recognition can be categorized into basic and compound types. Face attack detection includes deepfake detection and face anti-spoofing. Face recognition involves five scenarios: basic, cross-pose, cross-age, similar-looking, and occluded. Spatial relation understanding concerns relative position and count. Social relation understanding includes social relationship recognition and identity reasoning. Please refer to Appendix A.1 for detailed definitions and examples of these abilities.
@@ 2.2 Semi-Automatic Data Pipeline
Based on the hierarchical ability taxonomy defined in Section 2.1, we collect 16 public datasets from the face and human community, covering each L3 ability. Then, we employ a semi-automatic data pipeline to produce problems for the Face-Human-Bench.
@@
For ease of evaluation, we adopt multiple-choice as the problem format in our Face-Human-Bench. Each problem $P_{i}$ corresponds to a quadruple $(V_{i},Q_{i},O_{i},A_{i})$. Here, $V_{i}$ refers to the images obtained via the image processing pipeline $p_{image}:\mathbb{I}\rightarrow\mathbb{V}$. $p_{image}$ performs an operation such as cropping, concatenating, adding boxes, or leaving the original images unchanged, depending on the ability to test. $Q_{i}$ denotes the question. Each L3 ability includes a set of pre-written questions that share the same semantics but exhibit diversity. When producing samples, a question $Q_{i}$ is randomly selected from this question set. $O_{i}$ is the set of $n$ options $(o_{1},o_{2},...,o_{n})$, where $2\leq n\leq 4$. These options are obtained through the text processing pipeline $p_{text}:\mathbb{L}\rightarrow\mathbb{O}$. $p_{text}$ converts the original labels into one correct option and $n-1$ incorrect options. For some tasks, ChatGPT (OpenAI, 2023a) is used within $p_{text}$ to assist in generating incorrect options or adjusting options at the sentence level (fixing grammar or re-wording sentences for fluency. $A_{i}$ is the correct answer to the problem. The produced $P_{i}$ will be checked by data reviewers to ensure that the options are unambiguous and there is one and only one correct answer. The problems that do not meet the requirements will be removed.
In summary, our semi-automatic data pipeline leverages image and text processing pipelines, $p_{image}$ and $p_{text}$, to transform original samples into multiple-choice format problems. These problems are then manually checked to ensure quality. We obtain a benchmark with a development set of 900 problems for the MLLM community to evaluate during training iterations and a test set of 1800 problems for the formal evaluation in our paper. Additionally, the English problems are translated into Chinese to create a Chinese version of the benchmark. For more details on data sources, statistics, and the semi-automatic data pipeline, please refer to Sections A.2 and A.3.
@@ 3.1 Experimental Setup
Evaluation Protocols. We use the weighted accuracy of multiple-choice problems as the evaluation score. As shown in Figure 1, the proportion of the sectors represents the weight of the corresponding abilities in the overall score on the Face-Human-Bench. Note that we set equal weights for each L2 ability.  [^1] For detailed weights of each subset in Face-Human-Bench, please refer to Appendix A.2. To prevent models from favoring certain option letters over others, we shuffle the options to ensure the correct answers are evenly distributed across all option letters. During the testing, we add some constraint instructions to ensure MLLMs output only option letters as much as possible.  [^2] For the prompt template under zero-shot setting, please refer to Appendix B.2.1. After obtaining the MLLM’s response, we use regular expressions to extract the option letters. If this fails, we follow the implementation of MMBench (Liu et al., 2023c) using ChatGPT (OpenAI, 2023a) to extract the choices.  [^3] For the prompt for choice extraction, please refer to Appendix B.2.2.
Models. We evaluate 25 MLLMs in different sizes from 13 model families. For open-source models, we select LLaVA-13B (Liu et al., 2023b), LLaVA-1.5-7B/13B (Liu et al., 2023a), LLaVA-Next-7B/13B/34B (Liu et al., 2024a), MiniGPT-4-7B/13B (Zhu et al., 2024), InstructBLIP-7B/13B (Dai et al., 2023), Qwen-VL-Chat (Bai et al., 2023), InternLM-XComposer2-VL-7B (Dong et al., 2024), Yi-VL-6B (Young et al., 2024), InternVL-Chat-v1.2-Plus (Chen et al., 2023), InternVL-Chat-v1.5 (Chen et al., 2023), DeepSeek-VL-1.3B/7B-Chat (Lu et al., 2024), CogVLM2-19B-Chat (Hong et al., 2024), GLM-4V-9B (Hong et al., 2024), LLaVA-OneVison-0.5B/7B (Li et al., 2024). For closed-source models, we use Gemini-1.5-Pro (Reid et al., 2024), Claude-3.5-Sonnet (Anthropic, 2024a), GPT-4V (OpenAI, 2023b), and GPT-4o OpenAI (2024). For more details on these models, please refer to Appendix B.1.
Table 1: Zero-shot scores of MLLMs on the hierarchical Face-Human-Bench (EN). The highest scores for open-source and closed-source MLLMs are marked in blue and green respectively.
@@
Yi-VL-6B 75.5 51.7 65.0 52.0 58.5 34.0 43.0 38.5 50.0 48.0 48.0 50.0 44.0 48.0
MiniGPT-4-7B 24.0 17.7 26.0 24.0 25.0 31.5 40.5 36.0 38.0 56.0 44.0 48.0 34.0 44.0
InstructBLIP-7B 39.5 36.7 38.0 40.0 39.0 50.5 53.0 51.8 52.0 58.0 48.0 52.0 54.0 52.8
Qwen-VL-Chat 55.5 49.7 65.0 50.0 57.5 51.0 54.0 52.5 66.0 52.0 54.0 58.0 54.0 56.8
@@
LLaVA-1.5-7B 61.0 49.3 62.0 58.0 60.0 55.5 55.0 55.3 54.0 52.0 50.0 56.0 50.0 52.4
LLaVA-NeXT-7B 69.5 50.0 72.0 62.0 67.0 59.5 58.5 59.0 62.0 50.0 48.0 56.0 50.0 53.2
@@
CogVLM2-19B-Chat 75.0 57.3 71.0 70.0 70.5 37.0 51.0 44.0 66.0 36.0 44.0 46.0 48.0 48.0
GLM-4V-9B 79.5 55.7 79.0 74.0 76.5 46.0 50.0 48.0 68.0 54.0 54.0 62.0 52.0 58.0
MiniGPT-4-13B 20.5 24.3 35.0 26.0 30.5 49.5 37.5 43.5 52.0 46.0 42.0 46.0 48.0 46.8
InstructBLIP-13B 25.5 38.3 50.0 42.0 46.0 57.5 52.0 54.8 48.0 52.0 52.0 50.0 52.0 50.8
LLaVA-13B 32.0 40.7 56.0 30.0 43.0 55.0 54.0 54.5 52.0 60.0 52.0 40.0 52.0 51.2
LLaVA-1.5-13B 75.5 58.7 72.0 54.0 63.0 51.0 54.0 52.5 54.0 48.0 54.0 48.0 50.0 50.8
LLaVA-NeXT-13B 77.5 46.7 71.0 52.0 61.5 50.0 54.0 52.0 58.0 54.0 54.0 56.0 56.0 55.6
InternVL-Chat-v1.5 92.0 61.7 72.0 68.0 70.0 71.5 67.0 69.2 90.0 60.0 60.0 60.0 52.0 64.4
LLaVA-NeXT-34B 95.0 58.7 80.0 62.0 71.0 63.5 60.5 62.0 92.0 70.0 70.0 72.0 56.0 72.0
@@
Gemini-1.5-Pro 66.0 40.0 72.0 48.0 60.0 31.0 21.0 26.0 98.0 82.0 86.0 90.0 72.0 85.6
Claude-3.5-Sonnet 83.5 54.0 73.0 32.0 52.5 55.0 45.0 50.0 92.0 64.0 76.0 74.0 66.0 74.4
GPT-4V 77.5 53.7 75.0 48.0 61.5 50.5 58.5 54.5 96.0 72.0 92.0 82.0 64.0 81.2
GPT-4o 77.0 61.0 83.0 62.0 72.5 53.0 64.0 58.5 96.0 72.0 74.0 76.0 50.0 73.6
@@
Model Attr. Action RPU CC Mean SRR IR Mean Re-ID Face Human Per. Rea. Overall
@@
Yi-VL-6B 67.0 73.0 54.0 24.0 39.0 48.0 66.0 57.0 47.0 54.4 56.6 60.7 47.8 55.5
MiniGPT-4-7B 15.5 27.0 18.0 16.7 17.3 24.0 34.0 29.0 44.0 29.3 26.6 24.2 33.6 27.9
InstructBLIP-7B 31.0 46.0 34.0 0.7 17.3 16.0 28.0 22.0 51.0 43.9 33.5 40.7 35.8 38.7
Qwen-VL-Chat 49.5 83.0 54.0 34.0 44.0 64.0 70.0 67.0 50.0 54.4 58.7 57.9 54.5 56.5
@@
LLaVA-1.5-7B 62.0 71.0 54.0 30.0 42.0 68.0 78.0 73.0 63.0 55.6 62.2 59.8 57.6 58.9
LLaVA-NeXT-7B 62.0 80.0 62.0 24.7 43.3 62.0 86.0 74.0 56.0 59.7 63.1 64.6 56.6 61.4
@@
CogVLM2-19B-Chat 70.5 93.0 68.0 33.3 50.7 74.0 92.0 83.0 56.0 59.0 70.6 68.4 59.4 64.8
GLM-4V-9B 85.5 94.0 62.0 32.0 47.0 68.0 88.0 78.0 67.0 63.5 74.3 73.2 62.5 68.9
MiniGPT-4-13B 19.5 46.0 42.0 17.3 29.7 30.0 50.0 40.0 48.0 33.1 36.6 30.7 41.1 34.9
InstructBLIP-13B 33.5 71.0 38.0 28.0 33.0 52.0 86.0 69.0 51.0 43.1 51.5 44.9 51.0 47.3
LLaVA-13B 27.0 66.0 36.0 30.7 33.3 38.0 76.0 57.0 55.0 44.3 47.7 43.9 49.1 46.0
LLaVA-1.5-13B 60.5 72.0 44.0 26.0 35.0 60.0 60.0 60.0 54.0 60.1 56.3 63.7 50.0 58.2
LLaVA-NeXT-13B 69.5 74.0 46.0 28.0 37.0 58.0 70.0 64.0 63.0 58.7 61.5 63.5 54.9 60.1
InternVL-Chat-v1.5 89.5 89.0 62.0 50.7 56.3 70.0 74.0 72.0 77.0 71.5 76.8 78.6 67.4 74.1
LLaVA-NeXT-34B 91.5 88.0 64.0 59.3 61.7 64.0 86.0 75.0 88.0 71.7 80.8 77.7 74.2 76.3

### v3

@@ Abstract
Faces and humans are crucial elements in social interaction and are widely included in everyday photos and videos. Therefore, a deep understanding of faces and humans will enable multi-modal assistants to achieve improved response quality and broadened application scope. Currently, the multi-modal assistant community lacks a comprehensive and scientific evaluation of face and human understanding abilities. In this paper, we first propose a hierarchical ability taxonomy that includes three levels of abilities. Then, based on this taxonomy, we collect images and annotations from publicly available datasets in the face and human community and build a semi-automatic data pipeline to produce problems for the new benchmark. Finally, the obtained Face-Human-Bench includes a development set and a test set, each with 1800 problems, supporting both English and Chinese. We conduct evaluations over 25 mainstream multi-modal large language models (MLLMs) with our Face-Human-Bench, focusing on the correlation between abilities, the impact of the relative position of targets on performance, and the impact of Chain of Thought (CoT) prompting on performance. We also explore which abilities of MLLMs need to be supplemented by specialist models. The dataset and evaluation code have been made publicly available at https://face-human-bench.github.io/.
@@ 1 Introduction
Faces and humans are always the most crucial elements of photos and videos in our everyday lives. Consequently, they are also critical focuses in multi-modal AI applications. In the past two years, ChatGPT ChatGPT and GPT-4 GPT-4V have achieved great success with impressive instruction-following and multi-modal understanding capabilities respectively. Numerous excellent works LLaVA ; MiniGPT-4 ; InstructBLIP ; Qwen-VL from the open-source community have followed, collectively presenting the immense potential of multi-modal assistants. Since faces and humans are central to social interaction, a deep understanding of this information can make multi-modal assistants achieve improved response quality and broadened application scope. For instance, in movie understanding Movie101 ; AutoAD ; C-AD , identifying characters is a prerequisite for multi-modal assistants to describe the plot accurately. In multi-modal human-computer interaction VITA , perceiving expressions and body language can help multi-modal assistants accurately understand the context, generating more personalized and humanized responses. In media forensics FKA-Owl ; DFA-GPT ; CSR-DFD , determining whether deepfake artifacts exist on a face is crucial for multi-modal assistants to detect misinformation.
@@
Figure 1: The three-level ability taxonomy for evaluating face and human understanding abilities. We construct the Face-Human-Bench based on this taxonomy. The proportion of the sectors represents the weight of the corresponding abilities in the overall score on the Face-Human-Bench.
Comprehensive and scientific evaluation is the foundation for researching applications of multi-modal assistants related to “faces and humans." Existing benchmarks MME ; SEED-Bench ; MMbench for large multi-modal models typically involve limited abilities of face and human understanding, such as celebrity recognition, action recognition, identity reasoning, and social relation, leaving many important abilities unexplored. On the other hand, since face and human understanding is one of the earliest research topics in artificial intelligence, there are numerous datasets available for evaluating the performance of specialist models. The images and annotations from these datasets can serve as original material to evaluate multi-modal assistants.
As the starting point of our evaluation, we propose a hierarchical ability taxonomy, as shown in Figure˜1. This taxonomy consists of three levels. Level-1 (L1) has two perspectives to study: from the target perspective, L1 includes face understanding and human understanding; from the cognitive process perspective, L1 includes perception and reasoning. Subsequently, we incorporate finer-grained abilities into the taxonomy and categorize them into 10 Level-2 (L2) and 18 Level-3 (L3) ability dimensions. Then, based on this taxonomy, we collect datasets from the face and human community and use a semi-automatic data pipeline to transform original images and annotations into multi-modal QAs. The final obtained benchmark called Face-Human-Bench, including a development set and a test set, each with 1800 problems, supporting evaluations in both English and Chinese. For ease of evaluation, we adopt multiple-choice as the problem format following MMBench MMbench and SEED-Bench SEED-Bench .
Our study aims to provide readers with the following insights: Q1: How do existing Multi-modal Large Language Models (MLLMs) perform in the face and human understanding? Specifically, we focus on (a) the performance of 25 mainstream MLLMs, (b) the correlation between abilities at different levels, (c) the impact of the relative position of targets on performance, and (d) the impact of Chain of Thought (CoT) Multimodal-CoT prompting on performance. Q2: In the field of face and human understanding, which tasks’ specialized models achieve significantly better performance than current MLLMs? With these insights, researchers in the multi-modal community can strategically strengthen the deficient abilities in their general-purpose models. Moreover, researchers in the face and human understanding community can use our evaluation results to select suitable MLLMs as initialization for downstream application scenarios. Knowing the specific abilities that MLLMs lack, researchers can also strategically leverage the outputs of specialized models to construct multi-modal agents Visual-ChatGPT ; MM-REACT that enhance responses.
@@
In response to Q1, our main findings are as follows: (a) The Face-Human-Bench can effectively distinguish the abilities of MLLMs in faces and human understanding. Under the zero-shot setting, the best-performing closed-source model, GPT-4o GPT-4o , does not perform as well as the best open-source model, InternVL-Chat-v1.2-Plus InternVL . (b) The correlation coefficients can reveal correlations between abilities at different levels. At L2 and L3, there are some ability groups in which the ability dimensions exhibit significant positive correlations between each pair. (c) Many models show substantial performance differences on the same task with different relative positions of targets. We use a new metric called the relative position sensitivity score (RPSS) to measure this phenomenon. On this metric, InternLM-XComposer2-VL-7B Internlm-xcomposer2 performs the best, indicating that its performance is almost unaffected by the relative position of targets. (d) Introducing hints and CoT instructions into the prompts significantly improves the performance of the closed-source model GPT-4o, but has no effect on the open-source model, InternVL-Chat-v1.2-Plus. In response to Q2, we find that in tasks of deepfake detection, crowd counting, and face recognition (under challenging scenarios), the performance of MLLMs is significantly inferior to that of corresponding specialist models.
@@
  • We propose the Face-Human-Bench, the first benchmark dedicated to evaluating multi-modal assistants’ face and human understanding abilities. It is based on a three-level ability taxonomy and supports both English and Chinese.
@@
  • We explore which face and human understanding tasks see specialist models significantly outperform MLLMs, providing recommendations for downstream application scenarios.
@@ 2.1 Hierarchical Ability Taxonomy
As shown in Figure˜1, the proposed ability taxonomy includes three levels. Level 1 (L1) has two research perspectives. From the target perspective, L1 includes face understanding and human understanding. From the cognitive process perspective, L1 includes perception and reasoning. In our evaluation, perception involves direct comprehension of only one target, while reasoning requires synthesizing information from multiple targets and environments to conclude. There are ten abilities in total at Level 2 (L2). Five are focused on faces: facial attribute recognition, age estimation, facial expression recognition, face attack detection, and face recognition, and five are focused on humans: human attribute recognition, action recognition, spatial relation understanding, social relation understanding, and person re-identification. It should be noted that at L2, there are 6 abilities under perception and 4 abilities under reasoning. Level 3 (L3) further refines the ability dimensions at L2. Facial expression recognition can be categorized into basic and compound types. Face attack detection includes deepfake detection and face anti-spoofing. Face recognition involves five scenarios: basic, cross-pose, cross-age, similar-looking, and occluded. Spatial relation understanding concerns relative position and count. Social relation understanding includes social relationship recognition and identity reasoning. Please refer to Section˜A.1 for detailed definitions and examples of these abilities.
@@ 2.2 Semi-Automatic Data Pipeline
Based on the hierarchical ability taxonomy defined in Section˜2.1, we collect 16 public datasets from the face and human community, covering each L3 ability. Then, we employ a semi-automatic data pipeline to produce problems for the Face-Human-Bench.
@@
For ease of evaluation, we adopt multiple-choice as the problem format in our Face-Human-Bench. Each problem $P_{i}$ corresponds to a quadruple $(V_{i},Q_{i},O_{i},A_{i})$. Here, $V_{i}$ refers to the images obtained via the image processing pipeline $p_{image}:\mathbb{I}\rightarrow\mathbb{V}$. $p_{image}$ performs an operation such as cropping, concatenating, adding boxes, or leaving the original images unchanged, depending on the ability to test. $Q_{i}$ denotes the question. Each L3 ability includes a set of pre-written questions that share the same semantics but exhibit diversity. When producing samples, a question $Q_{i}$ is randomly selected from this question set. $O_{i}$ is the set of $n$ options $(o_{1},o_{2},...,o_{n})$, where $2\leq n\leq 4$. These options are obtained through the text processing pipeline $p_{text}:\mathbb{L}\rightarrow\mathbb{O}$. $p_{text}$ converts the original labels into one correct option and n1n-1 incorrect options. For some tasks, ChatGPT ChatGPT is used within $p_{text}$ to assist in generating incorrect options or adjusting options at the sentence level (fixing grammar or re-wording sentences for fluency). $A_{i}$ is the correct answer to the problem. The produced $P_{i}$ will be checked by data reviewers to ensure that options are unambiguous and there is one and only one correct answer. Problems that do not meet the requirements will be removed.
In summary, our semi-automatic data pipeline leverages image and text processing pipelines, $p_{image}$ and $p_{text}$, to transform original samples into multiple-choice format problems. These problems are then manually checked to ensure quality. We obtain a benchmark with a development set of 1800 problems for the MLLM community to evaluate during training and a test set of 1800 problems for the formal evaluation in our paper. Additionally, the English problems are translated into Chinese to create a Chinese version of the benchmark. For more details on data sources, statistics, and the semi-automatic data pipeline, please refer to Sections˜A.2 and A.3.
@@ 3.1 Experimental Setup
We use the weighted accuracy of multiple-choice problems as the evaluation score. As shown in Figure˜1, the proportion of the sectors represents the weight of the corresponding abilities in the overall score on the Face-Human-Bench. Note that we set equal weights for each L2 ability. [^1] For detailed weights of each subset, please refer to Section A.2. To prevent models from favoring certain option letters over others, we shuffle the options to ensure the correct answers are evenly distributed across all option letters. During the testing, we add some constraint instructions to ensure MLLMs output only option letters as much as possible. [^2] For the prompt template, please refer to Section B.2.1. After obtaining the MLLM’s response, we use regular expressions to extract the option letters. If this fails, we follow the implementation of MMBench MMbench using ChatGPT ChatGPT to extract the choices. [^3] For the prompt for choice extraction, please refer to Section B.2.2. We evaluate 25 MLLMs (as shown in Table˜1) in different sizes from 13 model families. For more details on these models, please refer to Section˜B.1.
Table 1: Zero-shot scores of MLLMs on the hierarchical Face-Human-Bench (EN). The highest scores for open-source and closed-source MLLMs are marked in blue and green respectively. The scores in the “random" row are theoretical values. For convenience, the names of the various tasks are abbreviated, where the abbreviations are formed by retaining the initial letters of phrases or the first few letters of key terms; their full names can be referred to in Figure˜1.
@@
Yi-VL-6B Yi-VL 75.5 51.7 65.0 52.0 58.5 34.0 43.0 38.5 50.0 48.0 48.0 50.0 44.0 48.0
MiniGPT-4-7B MiniGPT-4 24.0 17.7 26.0 24.0 25.0 31.5 40.5 36.0 38.0 56.0 44.0 48.0 34.0 44.0
InstructBLIP-7B InstructBLIP 39.5 36.7 38.0 40.0 39.0 50.5 53.0 51.8 52.0 58.0 48.0 52.0 54.0 52.8
Qwen-VL-Chat Qwen-VL 55.5 49.7 65.0 50.0 57.5 51.0 54.0 52.5 66.0 52.0 54.0 58.0 54.0 56.8
@@
LLaVA-1.5-7B LLaVA-1.5 61.0 49.3 62.0 58.0 60.0 55.5 55.0 55.3 54.0 52.0 50.0 56.0 50.0 52.4
LLaVA-NeXT-7B LLaVA-Next 69.5 50.0 72.0 62.0 67.0 59.5 58.5 59.0 62.0 50.0 48.0 56.0 50.0 53.2
@@
CogVLM2-19B-Chat CogVLM2 75.0 57.3 71.0 70.0 70.5 37.0 51.0 44.0 66.0 36.0 44.0 46.0 48.0 48.0
GLM-4V-9B CogVLM2 79.5 55.7 79.0 74.0 76.5 46.0 50.0 48.0 68.0 54.0 54.0 62.0 52.0 58.0
MiniGPT-4-13B MiniGPT-4 20.5 24.3 35.0 26.0 30.5 49.5 37.5 43.5 52.0 46.0 42.0 46.0 48.0 46.8
InstructBLIP-13B InstructBLIP 25.5 38.3 50.0 42.0 46.0 57.5 52.0 54.8 48.0 52.0 52.0 50.0 52.0 50.8
LLaVA-13B LLaVA 32.0 40.7 56.0 30.0 43.0 55.0 54.0 54.5 52.0 60.0 52.0 40.0 52.0 51.2
LLaVA-1.5-13B LLaVA-1.5 75.5 58.7 72.0 54.0 63.0 51.0 54.0 52.5 54.0 48.0 54.0 48.0 50.0 50.8
LLaVA-NeXT-13B LLaVA-Next 77.5 46.7 71.0 52.0 61.5 50.0 54.0 52.0 58.0 54.0 54.0 56.0 56.0 55.6
InternVL-Chat-v1.5 InternVL 92.0 61.7 72.0 68.0 70.0 71.5 67.0 69.2 90.0 60.0 60.0 60.0 52.0 64.4
LLaVA-NeXT-34B LLaVA-Next 95.0 58.7 80.0 62.0 71.0 63.5 60.5 62.0 92.0 70.0 70.0 72.0 56.0 72.0
@@
Gemini-1.5-Pro Gemini-1.5 66.0 40.0 72.0 48.0 60.0 31.0 21.0 26.0 98.0 82.0 86.0 90.0 72.0 85.6
Claude-3.5-Sonnet Claude-3.5 83.5 54.0 73.0 32.0 52.5 55.0 45.0 50.0 92.0 64.0 76.0 74.0 66.0 74.4
GPT-4V GPT-4V 77.5 53.7 75.0 48.0 61.5 50.5 58.5 54.5 96.0 72.0 92.0 82.0 64.0 81.2
GPT-4o GPT-4o 77.0 61.0 83.0 62.0 72.5 53.0 64.0 58.5 96.0 72.0 74.0 76.0 50.0 73.6
@@
Model Attr. Action RPU $C$ Mean SRR IR Mean Re-ID Face Human Per. Rea. Overall
@@
Yi-VL-6B Yi-VL 67.0 73.0 54.0 24.0 39.0 48.0 66.0 57.0 47.0 54.4 56.6 60.7 47.8 55.5
MiniGPT-4-7B MiniGPT-4 15.5 27.0 18.0 16.7 17.3 24.0 34.0 29.0 44.0 29.3 26.6 24.2 33.6 27.9
InstructBLIP-7B InstructBLIP 31.0 46.0 34.0 0.7 17.3 16.0 28.0 22.0 51.0 43.9 33.5 40.7 35.8 38.7
Qwen-VL-Chat Qwen-VL 49.5 83.0 54.0 34.0 44.0 64.0 70.0 67.0 50.0 54.4 58.7 57.9 54.5 56.5
@@
LLaVA-1.5-7B LLaVA-1.5 62.0 71.0 54.0 30.0 42.0 68.0 78.0 73.0 63.0 55.6 62.2 59.8 57.6 58.9
LLaVA-NeXT-7B LLaVA-Next 62.0 80.0 62.0 24.7 43.3 62.0 86.0 74.0 56.0 59.7 63.1 64.6 56.6 61.4
@@
CogVLM2-19B-Chat CogVLM2 70.5 93.0 68.0 33.3 50.7 74.0 92.0 83.0 56.0 59.0 70.6 68.4 59.4 64.8
GLM-4V-9B CogVLM2 85.5 94.0 62.0 32.0 47.0 68.0 88.0 78.0 67.0 63.5 74.3 73.2 62.5 68.9
MiniGPT-4-13B MiniGPT-4 19.5 46.0 42.0 17.3 29.7 30.0 50.0 40.0 48.0 33.1 36.6 30.7 41.1 34.9
InstructBLIP-13B InstructBLIP 33.5 71.0 38.0 28.0 33.0 52.0 86.0 69.0 51.0 43.1 51.5 44.9 51.0 47.3
LLaVA-13B LLaVA 27.0 66.0 36.0 30.7 33.3 38.0 76.0 57.0 55.0 44.3 47.7 43.9 49.1 46.0
LLaVA-1.5-13B LLaVA-1.5 60.5 72.0 44.0 26.0 35.0 60.0 60.0 60.0 54.0 60.1 56.3 63.7 50.0 58.2
LLaVA-NeXT-13B LLaVA-Next 69.5 74.0 46.0 28.0 37.0 58.0 70.0 64.0 63.0 58.7 61.5 63.5 54.9 60.1
InternVL-Chat-v1.5 InternVL 89.5 89.0 62.0 50.7 56.3 70.0 74.0 72.0 77.0 71.5 76.8 78.6 67.4 74.1
LLaVA-NeXT-34B LLaVA-Next 91.5 88.0 64.0 59.3 61.7 64.0 86.0 75.0 88.0 71.7 80.8 77.7 74.2 76.3